# Reinforcement Learning-based Receding Horizon Control using Adaptive Control Barrier Functions for Safety-Critical Systems

### v1

@@ Abstract
Optimal control methods provide solutions to safety-critical problems but easily become intractable. Control Barrier Functions (CBFs) have emerged as a popular technique that facilitates their solution by provably guaranteeing safety, through their forward invariance property, at the expense of some performance loss. This approach involves defining a performance objective alongside CBF-based safety constraints that must always be enforced. Unfortunately, both performance and solution feasibility can be significantly impacted by two key factors: (i) the selection of the cost function and associated parameters, and (ii) the calibration of parameters within the CBF-based constraints, which capture the trade-off between performance and conservativeness. To address these challenges, we propose a Reinforcement Learning (RL)-based Receding Horizon Control (RHC) approach leveraging Model Predictive Control (MPC) with CBFs (MPC-CBF). In particular, we parameterize our controller and use bilevel optimization, where RL is used to learn the optimal parameters while MPC computes the optimal control input. We validate our method by applying it to the challenging automated merging control problem for Connected and Automated Vehicles (CAVs) at conflicting roadways. Results demonstrate improved performance and a significant reduction in the number of infeasible cases compared to traditional heuristic approaches used for tuning CBF-based controllers, showcasing the effectiveness of the proposed method.
@@ I Introduction
The QP-CBF method has been applied to safety-critical control problems involving Connected and Automated Vehicles (CAVs) at various conflict points of transportation networks [12, 13, 14]. Aside from the issues addressed above, the solutions to these problems have also been largely limited to the longitudinal dynamics of vehicles, decoupling these from their lateral dynamics. In this context, we make the following contributions:
@@ V Multi-Agent Control of CAVs
where, $r_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}},\ \mu_{i}=\tan^{-1}\left(\frac{y_{i}}{x_{i}}\right)$; similarly, $r_{i_{c}}=\sqrt{x_{i_{c}}^{2}+y_{i_{c}}^{2}},\mu_{i_{c}}=\tan^{-1}\left(\frac{y_{i_{c}}}{x_{i_{c}}}\right)$. Here $\Phi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ may be any continuously differentiable function as long as it is strictly increasing and satisfies the boundary conditions $\Phi(x_{i}(t_{i}^{0}),y_{i}(t_{i}^{0}))=0$ and $\Phi(x_{i}(t_{i}^{m}),y_{i}(t_{i}^{m}))=\varphi$ [22]. Here $\varphi$ denotes the reaction time (as a rule, $\varphi=1.8s$ is used, e.g., [23]) and $\delta$ is a given minimum safe distance.

### v2

# Christos G. Cassandras2, Ioannis Ch. Paschalidis3, Wenchao Li2

@@ Abstract
Optimal control methods provide solutions to safety-critical problems but easily become intractable. Control Barrier Functions (CBFs) have emerged as a popular technique that facilitates their solution by provably guaranteeing safety, through their forward invariance property, at the expense of some performance loss.This approach involves defining a performance objective alongside CBF-based safety constraints that must always be enforced. Unfortunately, both performance and solution feasibility can be significantly impacted by two key factors such as the selection of the cost function and associated parameters, and the calibration of parameters within the CBF-based constraints, which capture the trade-off between performance and conservativeness. To address these challenges, we propose a Reinforcement Learning (RL)-based Receding Horizon Control (RHC) approach leveraging Model Predictive Control (MPC) with CBFs (MPC-CBF). In particular, we parameterize our controller and use bilevel optimization, where RL is used to learn the optimal parameters while MPC computes the optimal control input. We validate our method by applying it to the challenging automated merging control problem for Connected and Automated Vehicles (CAVs) at conflicting roadways. Results demonstrate improved performance and a significant reduction in the number of infeasible cases compared to traditional heuristic approaches used for tuning CBF-based controllers, showcasing the effectiveness of the proposed method.
@@ I Introduction
The QP-CBF method has been applied to safety-critical control problems involving Connected and Automated Vehicles (CAVs) at various conflict points of transportation networks [12, 13, 14, 15]. Aside from the issues addressed above, the solutions to these problems have also been largely limited to the longitudinal dynamics of vehicles, decoupling these from their lateral dynamics. In this context, we make the following contributions:
@@ V Multi-Agent Control of CAVs
where, $r_{i}=\sqrt{x_{i}^{2}+y_{i}^{2}},\ \mu_{i}=\tan^{-1}\left(\frac{y_{i}}{x_{i}}\right)$; similarly, $r_{i_{c}}=\sqrt{x_{i_{c}}^{2}+y_{i_{c}}^{2}},\mu_{i_{c}}=\tan^{-1}\left(\frac{y_{i_{c}}}{x_{i_{c}}}\right)$. Here $\Phi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ may be any continuously differentiable function as long as it is strictly increasing and satisfies the boundary conditions $\Phi(x_{i}(t_{i}^{0}),y_{i}(t_{i}^{0}))=0$ and $\Phi(x_{i}(t_{i}^{m}),y_{i}(t_{i}^{m}))=\varphi$. Here $\varphi$ denotes the reaction time (as a rule, $\varphi=1.8s$ is used, e.g., [24]) and $\delta$ is a given minimum safe distance, (see [25] for details).